Suppose that the monthly cost of a long-distance phone plan (in dollars) is a linear function of the total calling time (in minutes). When graphed, the function gives a line with a slope of 0.09.
The monthly cost for 35 minutes of calls is $15.94. What is the monthly cost for 29 minutes of calls?

Answer by ankor@dixie-net.com(22740)

Suppose that the monthly cost of a long-distance phone plan (in dollars) is a linear function of the total calling time (in minutes). When graphed, the function gives a line with a slope of 0.09.
The monthly cost for 35 minutes of calls is $15.94. What is the monthly cost for 29 minutes of calls?
:
write an equation using the point/slope formula y-y1 = m(x-x1)
from the problem; m = .09, x1 = 35, y1 = 15.94
:
y - 15.94 = .09(x - 35)
:
y - 15.94 = .09x - 3.15
:
y = .09x - 3.15 + 15.94
:
y = .09x + 12.79
:
:
What is the monthly cost for 29 minutes of calls?
substitute 29 for x in the above equation
y = .09(29) + 12.79
y = 2.61 + 12.79
y = $15.40 for 29 min